Prepare a statement of operations: (only 1 year)

Revenue from patients: $5,000,000, Medical services $ 600,000; Therapy services $100,000; support services $200,000, General services $ 300,000, Depreciation $150,000, Interest $ 50,000; interest income $1,000

------------------------------------------

Revenue:

Net Patient Service Revenue

Total Operating Revenue

Operating Expense:

....

Total Operating Expenses:

Income from operations:

Nonoperating Gains (Losses)

Interest Income:

Net Nonoperating Gains:

Revenue and Gains in Excess of Expenses and Losses:

Increase in Unrestricted Fund Balance:
Answer #1

Please see the table below. Please see the last column to understand what's going on.

Revenue: $ How
Net Patient Service Revenue 6,200,000 '=5000000+600000+100000+200000+300000 = A
Total Operating Revenue 6,200,000 B = A
Operating Expense:
Depreciation 150,000 C
Total Operating Expenses: 150,000 D = C
Income from operations: 6,050,000 E = B - D
Nonoperating Gains (Losses)
Interest Income: 1,000 F
Net Nonoperating Gains: 1,000 G = F
Revenue and Gains in Excess of Expenses and Losses: 6,051,000 H = E + G
Increase in Unrestricted Fund Balance: 6,051,000 '=H